OpenStudy (anonymous):

PLEASE HELP! Angle A in Quadrant III and 0degrees

OpenStudy (anonymous):

ok since tangent is positive in third Quad..so  using Pythagorean theorem \[\Large altitude=\sqrt{5^2-2^2}=\sqrt{16}=4\] so \[\Large tangent=\frac{altitude}{base}=\frac{4}{2}\]